Dav has received the following quote from a supplier of replacement ink cartridges used in laser jet printers:

Quantity

Unit Price

1

$25

2

$21

5

$19

10

$15

13

$14

Number of units/order

1

2

5

10

13

Price/unit (quoted) 0

$25

$21

$19

$15

$14

Total price / order 0

Price difference between orders

Quantity difference/order

Price/unit/order quantity difference

Assignment:

Using the format illustrated in Chapter 12, calculate a quantity discount analysis for a quote using specific quantities.

What quantity should Piet order if he wants to receive the lowest incremental costs between quantities?

Is there anything unusual about this quote?

Other than price and discounts, what must a buyer consider when determining what quantity to purchase?
Answer #1

1.In this discount approach based upon the quantity, once the quantity ordered is getting high, the discount rates also goes high. so a supplier with a bulk purchase will gain a lot as he'll order in bulk. So As you'll order more of the quantity, you'll be left with lesser price to pay .

2. If he wants to receive the lowest incremental cost he should order between 6 to 10 units as incremental cost is lowest =( 15*10)-(5*19)÷5

=55/5= 11 per unit

3. Yes, Incremental cost increases after 11 units to 13 units so one will buy lesser between 6 to 10 units.

4.A buyer should determine credit periods, credit terms, fiduciary relationship, after sales services, quality of products etc.